Q = – 3.8µC is placed at an origin. Determine (a) difference in potential and (b) electric field (magnitude and direction) between two points, A and B which are located at a distance of 62 cm north of the charge, and 88 cm west of the charge, respectively.
Answer #1

37 99 =-3-8MC =-3-8 x16°c, *A=0:620,1850.pem kW a) VA-VB - k OB VA VB = = KG (8B-VA NAVB giorno fe 0.88 m VA-VB = 93909x-3-86